What volume of oxygen gas in L at 5.6 atm and 27 °C would be needed to produce 187.6 grams of water? (10 pts) (First a stoichiometry problem, then a gas laws problem). Give your answer to 2 decimal places. 

2 H2 + O2 → 2H2O      


1
Expert's answer
2021-03-25T04:07:01-0400

Molar Mass of H2O = 18.01528

Molar Mass of O = 15.999

Ratio from the equation = 1 M of O = 2 M of H2O

187.6/18.01528 = 10.4134 moles of H2O

= 10.4134/ 2 = 5.2067 moles of O

1 M of O = 15.999

= 15.999 ×5.2067

= 83.3020

= 83.30 g or ml